Task 3: The company financials are given below: Net income €200m Depreciation expense €72m Interest expense...

Task 3:

The company financials are given below:

Net income

€200m

Depreciation expense

€72m

Interest expense

€120m

Increase in net working capital investment

€32m

Increase in fixed capital investment

€136m

Market value of debt

€1,440m

Cost of debt

8.5%

Cost of equity

14%

Tax rate

20%

Constant growth rate of free cash flow (forecasted)

4% (per year)

Capital Structure:

Debt

40%

Equity

60%

Number of shares outstanding

20,000,000

Find the following:

1.   Free cash flow to firm (actual),  Free cash flow to firm (next year forecast), Weighted average cost of capital, Value of the company (using FCF capitalization approach), Value of equity and Intrinsic value per share.
Answer #1

Equity Value = FCFE1 ke - gFirm Value = FCFFO x (1 + g) WACC - g FCFF1 WACC - 9Calculation of free cash flow to the firm Free Cash Flow to the Firm = = Net Income+ Non Cash Charges+ Interest(1-Tax Rate)-Value of Company Using FCF Capitalisation Approach Firm Value =200(1+4%)/(11.12-4)% *=200(1+4%)/7.12% =208/7.12% 2921 Firm